# VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

### v3

@@ 4 Experiments
We run two experiments with different settings for comparing VARCO Arena and current practice of reference-based benchmarking. In the first experiment (Section 4.2) we conduct a simulation study to control various factors that affect LLM benchmarking. This simulation tests our foundational propositions for VARCO Arena design (mentioned in Section 3.1 and 3.3) under a more controlled, simplified environment immune to noisy factors such as potential biases of LLM judges Park et al. (2024).
The other is empirical experiments (Section 4.3). We use gpt-4o and -mini as well as other several popular models such as Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to validate the effectiveness of our tournament approach against current LLM benchmarking practices. By presenting both simulation and empirical results, we aim to demonstrate the effectiveness of our tournament approach. In Section 4.1, we first describe the common experimental settings before getting into specific details of each experiments in the following subsections.
4.1 Chatbot Arena Leaderboard Ratings as Ground-Truth LLM Rankings
We compare the results of each benchmarking approach against the rankings from the Chatbot Arena leaderboard. Chatbot Arena is widely regarded as one of the most reliable leaderboards due to its extensive collection of human preference annotations. Given the large number of votes and the diverse set of prompts used for model comparisons, the resulting rankings are considered sufficiently accurate to serve as ground truth.
4.2 Experiment 1: Simulation Study
We designed a simple simulation to emulate a probabilistic model of LLM matches, adhering to the Elo rating system in a controlled environment. In line with the Elo model’s assumptions, our judge is configured to follow Equation 3 exactly. The judge stochastically determines the winner of each LLM match solely based on the Elo rating difference ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$). As described in Equation 3, the outcome of a single match is sampled according to $P_{\text{predict}}(i>j)$, which is computed as the product of the judge’s accuracy and the likelihood of model $i$ beating model $j$ based on the Elo gap ($P_{gt}$).
@@ 4.2 Experiment 1: Simulation Study
The details of our simulation settings are as follows:
Ground-truth Elo ratings (initial parameter of the simulation, and at the same time, the gold ranking to reproduce): We extracted Elo ratings from the English category of Chatbot Arena as of June 23. This Elo ratings are the estimates from massive user-submitted judgments (approximately 60% of the total submissions to the platform). For the simulation, any set of Elo ratings could be used, but we opted for real values computed from human preferences.
Judge Accuracy ($P_{\text{judge}}$): In practice, a judge’s accuracy is an adaptive value that depends on factors such as the specific prompt-response pair and the manner in which LLM judges are prompted. In this simulation study, we control this parameter, varying it from 0.6 to 0.9.
Number of Participant LLMs ($n_{\text{model}}$) and Benchmark Dataset Size ($|X|$): We varied these factors to assess the robustness of both the tournament and reference-based approaches under diverse conditions. This allowed us to explore how reliability changes with the number of participants in both data-poor and data-rich environments.
@@
    Select the participant LLMs and obtain their Elo ratings for the simulation.
@@
    Compute the expected relative win rates ($P_{\text{gt}}$, see Equation 3) from the participants’ Elo ratings.
@@
    Sample match outcomes for each LLM pair according to the benchmarking approach. The winner of each match is sampled from $P_{\text{pred}}$ (Equation 3), which depends solely on the Elo gap ($\Delta_{\text{ij}}$) and the judge’s accuracy ($P_{\text{judge}}$).
@@
    Repeat step 3 for the designated number of test prompts ($|X|$).
@@
    Calculate scores for ranking:
    1. (a)
      For the reference-based approach (current practice), use the win rate against the reference model (gpt-4-1106-preview, which has an Elo rating of 1233).
    2. (b)
      For the tournament approach (VARCO Arena), compute Elo ratings from all match outcomes.
@@
    Rank the models based on these scores.
@@
    Compute the Spearman correlation between the simulated rankings (from step 6) and the ground-truth rankings (from step 1).
We perform 50 trials for each simulation configuration to mitigate randomness from tournament brackets and sampling.
4.3 Experiment 2: Assessing VARCO Arena Using Various LLM Judges
To empirically validate our proposal, we evaluated the reliability of both VARCO Arena and reference-based approach over the top 20 models from the Chatbot Arena leaderboards. This experiment employs actual prompt inputs and LLM outputs, distinguishing it from the earlier simulation study.
@@ 4.3.1 Dataset: Test Prompts and LLM Responses Used
Testing the benchmarking approaches requires: (1) test prompts and (2) the corresponding responses from LLMs. For the benchmark dataset, we selected Arena-Hard-Auto Li et al. (2024). The prompts in Arena-Hard-Auto were carefully curated from Chatbot Arena user queries. This dataset consists of 500 prompts—two instances for each of 250 subtopics. Although AlpacaEval Li et al. (2023), which comprises 800 prompt-reference pairs, could serve as a viable testbed, we opted for Arena-Hard-Auto because its design aligns more closely with Chatbot Arena. Arena-Hard-Auto uses responses from gpt-4-0314 as the reference outputs. For ranking, we utilized the reserved outputs of the top 21 models from the Arena-Hard-Auto Browser. [^3] Extracted from the 2024 Jul 6 commit (fd42026).
@@ 4.3.2 Participant LLMs
For ranking, we selected 20 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
@@ 4.3.3 LLM Judges
We used several aligned LLMs as judges for testing both benchmarking approaches. LLMs of our choice are gpt-4o family of models OpenAI et al. (2024), Claude3.5, and a selection of open-weight models: Qwen2.5 Qwen et al. (2025), Llama3.1 Grattafiori et al. (2024), and Gemma2 Team et al. (2024). For pairwise comparisons of responses, we employed the judging prompt suggested in LLMBar Zeng et al. (2024) (See Appendix A.6.2). The same judge prompt was applied consistently across both the tournament and reference-based approaches. To mitigate position bias Wu and Aji (2023), the order of model responses was alternated during evaluation. Further details on the LLM-as-a-judge configuration are provided in Appendix A.6.
@@
Experiment 1 (Simulation Study): This experiment uses the ground truth Elo ratings of the models to initialize the simulation. We vary control parameters for the benchmarking approaches—including the judge’s accuracy ($P_{\text{judge}}$), the number of test prompts used ($|X|$), and the number of participant LLMs ($n_{\text{model}}$)—to determine which benchmarking approach more accurately reproduces the participants’ ranking. For each configuration, we conduct 50 trials of experiments.
Experiment 2 (Empirical Runs): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges. We select the top 20 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts. For both the tournament and reference-based approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings. In our empirical study, we conduct 500 trials for each experimental setting.
Figure 2: Simulation results comparing the tournament and reference-based approaches. The tournament method consistently outperforms the reference-based approach in Spearman correlation across various control variables: the number of participant LLMs ($n_{\text{models}}$), the number of benchmark prompts ($|X|$), and judge precision ($P_{\text{judge}}$).
@@ 5 Results and Discussion
We assess the reliability and robustness of VARCO Arena as a means for LLM benchmarking, comparing it against the current reference-based approach. Our results from both simulation study and empirical runs indicate that the tournament approach of VARCO Arena yields rankings that align more closely with the ground-truth Elo leaderboards. We present our findings using whisker plots and tables in the following sections.
5.1 Experiment 1: Simulation Study Results
Figure 2 illustrates noticeable differences in Spearman correlation, indicating that the tournament approach is more reliable than the reference-based method. The consistent performance gap across various conditions—namely, the number of participants, the number of test prompts, and judge accuracy ($n_{\text{model}}$, $|X|$, and $P_{\text{judge}}$)—demonstrates the robustness of the tournament approach. Although the simulation simplifies real-world complexity, a similar performance gap was observed in the empirical findings (Experiment 2, Figure 3). This consistency suggests that the robust performance of VARCO Arena is not coincidental or limited to a specific empirical setting of ours.
5.2 Experiment 2: Empirical Results
Figure 3: Results of VARCO Arena (tournament) and reference-based approach with gpt-4o (left) and gpt-4o-mini (right) judge. VARCO Arena constantly records higher Spearman correlation coherent with the Experiment 1 result (Figure 2). Results summary is on Table 2.
As hinted in the previous section, the empirical results in Figure 3 show that VARCO Arena consistently outperforms the reference-based approach. Although the performance gaps are less pronounced than in the simulation, the same trend persists. In Table 2, we report the median values for VARCO Arena and the reference-based approach using the gpt-4o family of judges while varying the number of test prompts ($|X|$). These results consistently demonstrate that VARCO Arena outperforms the reference-based method. Note that VARCO Arena shows similar or superior reliability even in extreme data-poor benchmark condition ($|X|=50$).
Table 3 presents the outcomes when using other LLMs as judges, with a fixed number of prompts ($|X|=500$). The results for Claude3.5-sonnet, Llama3.1-8b, and Qwen2.5-7b follow a similar trend. However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable for benchmarking. Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using VARCO Arena or considering reference-based approach.
@@ 5.2 Experiment 2: Empirical Results
comp. to ref. (4o) 0.895 0.935 0.963 0.966 0.964
tournament (4o) 0.905 0.940 0.960 0.970 0.970
comp. to ref. (4o-mini) 0.895 0.908 0.917 0.916 0.912
tournament (4o-mini) 0.901 0.919 0.931 0.933 0.933
Table 2: Spearman correlation ($\uparrow$) varying over size of the benchmark set ($|X|$) for each benchmarking approach. Comp. to ref. refers to reference-based approach.
$|X|$ = 500 claude3.5 sonnet llama3.1 8b-it qwen2.5 7b-it qwen2.5 0.5b-it gemma2 2b-it
comp. to ref. 0.924 0.820 0.756 0.089 0.592
tournament 0.930 0.850 0.811 -0.124 0.552
Table 3: Spearman correlation ($\uparrow$) result using other LLMs as a judge. Comp. to ref. refers to reference-based approach.
@@ 5.3 Incorporating a New LLM into an Existing Leaderboard
While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case for leaderboards. We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a reference. Our findings indicate that the latter approach is more reliable (Table 4). Further details and discussions are provided in Appendix A.4.
binary search (4o) 0.92 1.84 2.13 1.72
comp. to 1st (4o) 1.98 1.55 1.57 1.39
binary search (4o-mini) 1.27 1.82 1.21 1.5
comp. to 1st (4o-mini) 1.00 1.43 1.43 1.37
Table 4: Comparison of the binary search method versus using the top-performing model’s response as a reference (comp. to 1st) for inserting a new LLM into the leaderboard. We report the mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from the ground-truth leaderboard as an additional error metric. For further details, see Algorithm 2 in Appendix.
@@ 6 Related Works
6.1 Elo-based LLM Benchmarking
Recent studies have leveraged Elo ratings derived from human preferences to benchmark LLMs Boubdir et al. (2023). For instance, Chatbot Arena Chiang et al. (2024) employs Elo as an evaluation metric, while RAGElo Rackauckas et al. (2024) uses it as a ranking metric. These implementations highlight the benefits of applying Elo ratings for open-ended text quality evaluation.
6.2 Reference-free Evaluation
Reference-free evaluation has emerged as a promising alternative to static reference-based methods from the era of Natural Language Generation. Advances in LLM capabilities have enabled models to assess open-ended responses effectively Jauhiainen and Guerra (2024). When reference quality is poor, reference-free metrics like XComet Guerreiro et al. (2023) works as a more reliable alternatives.
@@ 7 Conclusion
We introduced VARCO Arena, a reference-free LLM benchmarking approach that employs a tournament-style framework with direct pairwise comparisons to evaluate the generative capabilities of LLMs. VARCO Arena offers a cost-effective, scalable, and adaptable solution for benchmarking response quality. Our simulation study and empirical experiments demonstrate that VARCO Arena consistently achieves higher rank reliability compared to current reference-based methods, as evidenced by stronger correlations with human-established Elo ratings. Given its robust performance and flexibility, we believe VARCO Arena can serve as a reliable automated tool for model selection and ranking across diverse and evolving use cases. Future work will explore broader applications, such as benchmarking LLMs in multi-modal settings (e.g., those incorporating visual or audio inputs and outputs).
@@
Although we tested robustness of the tournaments performed by VARCO Arena, it still has added factor for randomness such as match bracket initialization which does not applies to reference-based method. Rooms for improvement exist for more informative match-making algorithm that would achieve better ranking than single-elimination tournaments within same or less number of matches.

### v4

@@ 4 Experiments
We conducted two experiments to evaluate Arena-Lite against baseline-mediated benchmarking. The first experiment (Section 4.2) utilized a stochastic model to simulate LLM competitions, comparing Arena-Lite’s tournament-based direct comparison with baseline-mediated evaluation. This controlled setup allowed us to test Arena-Lite’s design principles, such as the effectiveness of direct versus mediated comparison (Section 3.1) and tournament-based sampling (3.3), while isolating variables and minimizing noise, such as LLM judge biases Park et al. (2024). The second experiment (Section 4.3) validates Arena-Lite empirically using various LLMs as judges and public benchmark data. We tested models including gpt-4o, gpt-4o-mini, Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to assess Arena-Lite’s effectiveness against standard benchmarking practices. Together, these experiments demonstrate the superior reliability and efficiency of Arena-Lite’s tournament approach. Section 4.1 outlines shared experimental settings, followed by detailed descriptions of each experiment in subsequent subsections.
4.1 Chatbot Arena Leaderboard as Ground-Truth Rankings
We benchmark Arena-Lite and baseline-mediated evaluation against rankings from the Chatbot Arena leaderboard, widely recognized for its reliability due to extensive human preference annotations. With a large volume of votes across diverse prompts, these rankings provide a robust ground truth for model comparisons.
4.2 Experiment 1: Controlled Stochastic Modeling of LLM Competitions
We suggest a simple stochastic model based on the Bradley-Terry (BT) framework to compare Arena-Lite’s approach with baseline-mediated evaluation. This experiments simulates prompt-agnostic LLM competitions which follows the Bradley-Terry model’s presumption, with outcomes determined by a judge following Equation 3. The judge’s decision is based on the BT preference difference ($\Delta_{\text{ij}}$) between models $i$ and $j$, and the judge’s accuracy ($P_{\text{judge}}$):
@@ 4.2 Experiment 1: Controlled Stochastic Modeling of LLM Competitions
With the model of judge above (Equation 3), we simulate both Arena-Lite’s tournament-based approach and baseline-mediated approaches according to the following initial conditions and procedures.
Initial conditions:
  • Ground-Truth BT Preference: We extracted BT preferences from the English category of Chatbot Arena (as of June 23), derived from approximately 60% of user-submitted judgments. These preferences serve as both the initial model parameters and the ground-truth rankings for evaluation.
  • Judge Accuracy ($P_{\text{judge}}$): We varied judge accuracy from 0.6 to 0.9.
[…]
@@
    Select participant LLMs and their BT preferences.
@@
    Compute expected win rates ($P_{\text{gt}}$) using Equation 3.
@@
    Sample match outcomes based on $P_{\text{predict}}$ (Equation 3), determined by the BT preference rating gap ($\Delta_{\text{ij}}$) and judge accuracy ($P_{\text{judge}}$).
@@
    Repeat for the specified number of test prompts ($|X|$).
@@
    Compute scores:
    • Baseline-mediated: Win rate against a reference model (gpt-4-1106-preview, rating 1233).
    • Arena-Lite: BT preference from all tournament match outcomes.
@@
    Rank models based on scores.
@@
    Calculate Spearman correlation between simulated and ground-truth rankings.
We conducted 50 trials per configuration to account for stochasticity in initial tournament brackets and judging process.
4.3 Experiment 2: Empirical Validation of Arena-Lite with real LLM Judge
To empirically validate our proposal, we evaluated the reliability of both Arena-Lite and baseline-mediated approach over the top 19 models from the Chatbot Arena leaderboards. This experiment employs actual prompt inputs and LLM outputs, distinguishing it from the earlier simulation study.
@@ 4.3.1 Dataset: Test Prompts and LLM Responses Used
Testing the benchmarking approaches requires: (1) test prompts and (2) the corresponding responses from LLMs. For the benchmark dataset, we selected Arena-Hard-Auto Li et al. (2024). The prompts in Arena-Hard-Auto were carefully curated from Chatbot Arena user queries. This dataset consists of 500 prompts—two instances for each of 250 subtopics. Although AlpacaEval Li et al. (2023), which comprises 800 prompt-reference pairs, could serve as a viable testbed, we opted for Arena-Hard-Auto because its design aligns more closely with Chatbot Arena. Arena-Hard-Auto uses responses from gpt-4-0314 as the baseline outputs. For ranking, we utilized the reserved outputs of the top 20 (=19 + baseline) models from the Arena-Hard-Auto Browser. [^2] Extracted from the 2024 Jul 6 commit (fd42026).
@@ 4.3.2 Participant LLMs
For ranking, we selected 19 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
@@ 4.3.3 LLM Judges
We used several aligned LLMs as judges for testing both benchmarking approaches. LLMs of our choice are gpt-4o family of models OpenAI et al. (2024), Claude3.5, and a selection of open-weight models: Qwen2.5 Qwen et al. (2025), Llama3.1 Grattafiori et al. (2024), and Gemma2 Team et al. (2024). For pairwise comparisons of responses, we employed the judging prompt suggested in LLMBar Zeng et al. (2024) (See Appendix A.8.2). The same judge prompt was applied consistently across both the tournament and baseline-mediated approaches. To mitigate position bias Wu and Aji (2023), the order of model responses was alternated during evaluation. Further details on the LLM-as-a-judge configuration are provided in Appendix A.8.
@@
[…]
Experiment 2 (Empirical Validation): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges. We select the top 19 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts. For both the tournament and baseline-mediated approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings. In our empirical study, we conduct 500 trials for each experimental setting.
Figure 2: Comparison of LLM ranking reliability between Arena-Lite and a baseline method in a stochastic simulation (Experiment 1, Sec. 4.2). Ranking reliability is measured by the Spearman correlation ($\uparrow$) between the competition-derived ranking and the ground-truth ranking. Each box plot summarizes the results from 50 trials. The subplots analyze the effect of varying (from left to right) the number of competing models ($n_{\text{models}}$), the number of prompts ($n_{\text{prompts}}$), and the accuracy of the judge ($P_{\text{judge}}$). The single-elimination structure of Arena-Lite results in consistently higher correlation scores.
@@ 5 Results and Discussion
We assess the reliability and robustness of Arena-Lite as a means for LLM benchmarking, comparing it against the current baseline-mediated approach. Our results from both simulation study and empirical runs indicate that the tournament approach of Arena-Lite yields rankings that align more closely with the ground-truth Chatbot Arena leaderboards. We present our findings using whisker plots and tables in the following sections.
5.1 Experiment 1: Modeling Experiment Results
[…]
5.2 Experiment 2: Empirical Validation Results
Figure 3: Ranking reliability of Arena-Lite vs. utilizing baseline outputs. Arena-Lite consistently demonstrates higher Spearman’s rank correlation across numbers of benchmark prompts ($|X|$), indicating more reliable ranking. The evaluation was performed using gpt-4o (left) and gpt-4o-mini (right) as judge models, with a fixed number of models ($n_{\text{models}}$=19). Each box plot summarizes the results of 50 runs. (Experiment 2, Sec. 4.3).
As hinted in the previous section, the empirical results in Figure 3 show that Arena-Lite consistently outperforms the baseline-mediated approach. Although the performance gaps are less pronounced than in the simulation, the same trend persists. In Table 2, we report the median values for Arena-Lite and the baseline-mediated approach using the gpt-4o family of judges while varying the number of test prompts ($|X|$). These results consistently demonstrate that Arena-Lite outperforms the baseline-mediated method. Note that Arena-Lite shows similar or superior reliability even in extreme data-poor benchmark condition ($|X|=50$).
Table 3 presents the outcomes when using other LLMs as judges, with a fixed number of prompts ($|X|=500$). The results for Claude3.5-sonnet, Llama3.1-8b, and Qwen2.5-7b follow a similar trend. However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable as an LLM judge. Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using Arena-Lite or considering baseline-mediated approach.
@@ 5.2 Experiment 2: Empirical Validation Results
baseline-mediated (4o) 0.895 0.935 0.963 0.966 0.964
Arena-Lite (4o) 0.905 0.940 0.960 0.970 0.970
baseline-mediated (4o-mini) 0.895 0.908 0.917 0.916 0.912
Arena-Lite (4o-mini) 0.901 0.919 0.931 0.933 0.933
Table 2: Robustness of ranking methods to benchmark set size, $|X|$ (Experiment 2, Sec. 4.3). The table shows the median Spearman correlation ($\uparrow$) from 500 trials. Arena-Lite consistently achieves higher correlation than the baseline-mediated approach across all dataset sizes ($|X|$), demonstrating its superior reliability and robustness for ranking LLMs.
$|X|$ = 500
claude3.5
sonnet
llama3.1
8b-it
qwen2.5
7b-it
qwen2.5
0.5b-it
gemma2
2b-it
baseline-mediated 0.924 0.820 0.756 0.089 0.592
Arena-Lite 0.930 0.850 0.811 -0.124 0.552
Table 3: Robustness of ranking methods to the choice of judge LLM (Experiment 2, Sec. 4.3). The table shows the Spearman correlation ($\uparrow$) between ground-truth LLM rankings and the results from each method. The values are medians from 500 trials. The results suggest that around 7B parameters-large LLMs is a viable minimum threshold for a reliable judge. Full results for other benchmark sizes are available in Appendix, Table 4.
@@ 5.3 Incorporating a New LLM into an Existing Leaderboard
While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case. We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a baseline. Our findings indicate that the later approach is more reliable (Table 5, Appendix). Further details and discussions are provided in Appendix A.6.
@@ 6 Related Works
6.1 LLM-as-a-Judge for Systems Ranking
Utilizing LLM-as-a-Judge as a building block for systems ranking has become a common practice in the LLM benchmarking community. Several studies have investigated how LLM judges compare to human evaluators, examining their similarities and differences Park et al. (2024), as well as how these differences impact system rankings (e.g., JuStRank Gera et al. (2024)Gao et al. (2025)). Our research extends these approaches by proposing a method that orchestrates LLM-as-a-Judge through a well-established tournament structure to derive rankings among systems.
6.2 Efficient and Reliable Evaluation
There is a growing body of research focused on optimizing the number of evaluations while maintaining reliability when using LLM-as-a-Judge for system ranking. Perlitz et al. proposed a metric called DIoR to quantify the relationship between computational costs and system ranking reliability. UniCBE Yuan et al. (2025) introduced a method to analyze the relationship between reliability and the number of judge evaluations based on uncertainty. BenchBench Perlitz et al. (2024b) systematically analyzed consistency across benchmarks and provided a package to facilitate this analysis. tinyBenchmarks Maia Polo et al. (2024) explored strategies to minimize the number of evaluations across various established benchmarks. Arena-Lite relates to these studies in that it leverages the properties of tournament structures and direct comparisons to achieve more reliable results with fewer judge evaluations.
@@ 7 Conclusion
We introduced Arena-Lite, an efficient and reliable framework for evaluating Large Language Models (LLMs) through tournament-based direct comparisons. By eliminating the need for baseline encore outputs and adopting head-to-head comparison, Arena-Lite achieves higher reliability in system rankings with reduced number of comparisons. Our experiments, encompassing controlled stochastic modeling and empirical validation with various LLM judges, confirm that Arena-Lite consistently outperforms standard baseline-mediated evaluation methods, even with smaller datasets or weaker judges. The release of an accessible web demonstration and code supports the adoption of Arena-Lite to help streamlining model development cycle across research and industry. Future work will extend Arena-Lite’s application to diverse domains, including multi-modal LLM evaluation involving visual or audio inputs and outputs.
@@
While we conducted extensive testing to assess the robustness of Arena-Lite tournaments—including 50 and 500 trials for Experiment 1 and Experiment 2, respectively—some inherent sources of randomness remain, such as variation due to initial match bracket assignments. The randomness in bracket assignment is added for adopting tournament structure of Arena-Lite and may influence outcome stability. Future work could explore more informative or adaptive matchmaking strategies that improve ranking fidelity beyond what is achievable with single-elimination formats, potentially within the same or even fewer number of matches.